OpenStudy (anonymous):

I still don't understand :( Solve this system of equations. 12x+8y=-52 6x+3y=-21

OpenStudy (anonymous):

Multiply both sides of the second equation by 2: 12x+8y=-52 12x+6y=-42 Now you can subtract the second equation from the first (this eliminates x). 2y=-10 Now solve for y: y=-5 To solve for x, input y=-5 into either of the original equations and solve: 6x+3(-5)=-21 This gives, 6x-15=-21 and, 6x=-6 Thus, x=-1 So you solutions are x=-1 and y=-5